# NotebookOS: A Notebook Operating System for Interactive Training with On-Demand GPUs

### v1

@@ 7. Conclusion
The key insight of this paper is that notebooks are long-running but with fragmented executions, rendering intermittent, sporadic, and often low GPU utilization. We built NotebookOS, a GPU-efficient notebook platform designed to meet the unique notebook workload requirements and prioritizes interactivity. NotebookOS oversubscribes and multiplexes GPUs using a novel replicated kernel design. NotebookOS dynamically allocates GPUs only for notebook cell executions that involve GPU training. We evaluated NotebookOS using production notebook workloads. Results show that NotebookOS reduces cluster GPU resource cost significantly compared to existing notebook platforms. Integrating NotebookOS in production Sensei is part of our future work.
Appendix A Appendix
A.1. Implementation
@@
A.2. Simulation Study: Resource Usage
@@
Figure 13. Simulated GPU usage timeline over a 90-day snippet of the SenseiTrace spanning June through August.
@@
A.3. Detailed Latency & Overhead Breakdown
Figure 14. Workflow steps of notebook cell execution requests in NotebookOS and other baselines.
Figure 15. Detailed end-to-end latency breakdown of execute requests messages processed by Reservation.
Figure 16. Detailed end-to-end latency breakdown of execute requests messages processed by Batch.
Figure 17. Detailed end-to-end latency breakdown of execute requests messages processed by NotebookOS.
Figure 18. Detailed end-to-end latency breakdown of execute requests messages processed by NotebookOS (LCP).
@@ A.3. Detailed Latency & Overhead Breakdown
A.4. Active Sessions & Training Events
Figure 19. The number of active user-submitted trainings and active user sessions during the full summer portion (i.e., June, July, and August) of the Sensei workload trace.
Figure 19 presents a timeline plot showing the number of active sessions during the “summer portion” of the Sensei workload trace. Specifically, the number of active sessions is plotted on the secondary (i.e., right) y-axis. The summer portion includes data from the entirety of the months of June, July, and August. Initially, there are 0 active user sessions. By the end of June, July, and August, there are 206, 312, and 397 active user sessions, respectively. For June, the mean and median number of active user sessions are 115 and 130, respectively. For July, the mean and median number of active user sessions are 233 and 234, respectively. Finally, for August, the mean and median number of active user sessions are 379 and 385, respectively. The maximum number of active sessions at any given time throughout the summer Sensei workload trace is 433.
Figure 19 also presents a timeline plot showing the number of user-submitted training events being processed during the “summer portion” of the Sensei workload trace. Specifically, the number of active sessions is plotted on the primary (i.e., left) y-axis. By the end of June, July, and August, there are 58, 83, and 73 active user-submitted trainings, respectively. The mean and median number of active user-submitted trainings for the entire summer are 67.6293 and 68, respectively. For June, the mean and median number of active user-submitted trainings are 31 and 30, respectively. For July, the mean and median number of active user-submitted trainings are 65 and 67, respectively. Finally, for August, the mean and median number of active user-submitted trainings are 105 and 108, respectively. The maximum number of active user-submitted trainings at any given time throughout the summer Sensei workload trace is 141.
A.5. Prototype Demonstration and Observability
Figure 20. NotebookOS’s admin dashboard and the main page of the workload orchestrator (implemented as a React app).
We have implemented an easy-to-use dashboard interface as part of the NotebookOS software to simplify deployment and reproducibility. Figure 20 shows a screenshot of NotebookOS’s admin dashboard for global observability.
(a) NotebookOS dashboard example 1.
(b) NotebookOS dashboard example 1.
Figure 21. Demonstration of NotebookOS’s Grafana dashboard displaying real-time metircs.
Figure 21 shows two screenshots of the Grafana NotebookOS dashboard displaying real-time metrics for NotebookOS’s system performance and server cluster resource usage. The dashboard includes various panels that visualize key performance indicators, such as CPU/memory/GPU utilization, subscription ratios, interactivity delays, and the number of active kernels. Additional panels track the number of hosts in the system, messages sent, and other critical operational metrics. The layout features time-series graphs, gauges, and tables, providing a comprehensive view of system health and efficiency at a glance.

### v2

@@ 8. Conclusion
The key insight of this paper is that notebooks are long-running but with fragmented executions, rendering intermittent, sporadic, and often low GPU utilization. We built NotebookOS, a GPU-efficient notebook platform designed to meet the unique notebook workload requirements and prioritizes interactivity. NotebookOS oversubscribes and multiplexes GPUs using a novel replicated kernel design. NotebookOS dynamically allocates GPUs only for notebook cell executions that involve GPU training. We evaluated NotebookOS using production notebook workloads. Results show that NotebookOS reduces cluster GPU resource cost significantly compared to existing notebook platforms. Integrating NotebookOS in production is part of our future work.
Acknowledgments
We thank the anonymous reviewers and our shepherd, Rodrigo Bruno, for their valuable feedback and comments, which improved the paper. This research was supported in part by NSF grants OAC-2411009 and CNS-2322860 (NSF CAREER Award). We also acknowledge support from NSF CloudBank for providing AWS credits, and thank Adobe for their generous research gift. Benjamin Carver was supported by a Presidential Scholarship from George Mason University.
Appendix A Artifact Appendix
A.1. Abstract
This section presents supplementary materials and guidance intended to support the reproducibility of NotebookOS and its experimental evaluation. Provided resources include the source code, dataset/model names, config files, and tools used in the experimental framework. Comprehensive instructions for deployment and carrying out the experiments are available within the linked GitHub repositories [78, 77].
A.2. Artifact Check-List (Meta-Information)
  • Model: VGG-16, ResNet-18, Inception v3, BERT, GPT-2, Deep Speech 2.
  • Data set: CIFAR-10, CIFAR-100, Tiny ImageNet, IMDb Large Movie Reviews, CoLA, LibriSpeech.
  • Run-time environment: AWS, Linux, Ubuntu, WSL 2.
  • Hardware: AWS EC2 virtual machines.
  • Metrics: Interactivity delay, task completion time, number of GPUs provisioned, number of GPUs actively used, profit, profit margin, GPU-hours saved.
  • Output: Numerical statistics.
  • Experiments: Execution of real-world workload trace excerpt on prototype implementation, extension of real-world experiments using simulation study.
  • How much disk space is required (approximately)?: 10’s of GB across multiple virtual machines.
  • How much time is needed to prepare workflow (approximately)?: Several hours.
  • How much time is needed to complete experiments (approximately)?: Several hours.
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: Pending.
  • Archived (provide DOI)?: 10.5281/zenodo.15832098.
A.3. Description
A.3.1. How to Access
NotebookOS can be deployed using a set of Ansible [3] playbooks and shell scripts that we have provided. These tools support deployment across an arbitrary number of AWS EC2 virtual machines, allowing users to customize the scale of their infrastructure. The latest versions of the Ansible playbooks are available in the NotebookOS GitHub repository.
A.3.2. Hardware and Software Dependencies
NotebookOS was developed, tested, and evaluated on WSL 2 (Windows Subsystem for Linux 2) and Ubuntu on AWS EC2. The WSL 2 environment used Ubuntu 22.04.5 LTS (Jammy) with Docker 27.2.0, Go 1.22.9 (linux/amd64), Python 3.12.6, Protoc 27.2, and CUDA 12.7 (via NVIDIA driver v566.36, NVML v565.77, and NVIDIA-SMI v565.77.01). The host OS was Windows 10 v22H2 (OS build 19045.5487), with WSL 2 kernel version 5.15.167.4-1 and WSLg version 1.0.65. The AWS EC2 environment ran Ubuntu 24.04.1 LTS (Noble) with Docker 27.3.1 (build ce12230) with matching versions of Golang, Python, and Protoc. Currently, the glibc version in the Global and Local Scheduler Docker containers is ldd (Debian GLIBC 2.36-9+deb12u7) 2.36; within the Jupyter Docker container, it is ldd (Ubuntu GLIBC 2.35-0ubuntu3.8) 2.35.
Remote Storage. NotebookOS persists intermediate workload data to remote storage and currently supports AWS S3 (recommended), Redis and HDFS. For S3, one must create an AWS IAM role with access to the necessary AWS S3 buckets. When provisioning the AWS EC2 virtual machines, be sure that they are assigned an AWS IAM role with access to the AWS S3 buckets intended for use by the distributed kernels.
A.3.3. Models and Datasets
To ensure a thorough evaluation of NotebookOS and the various other baselines, we integrated support for a variety of different deep learning models and datasets across several different application domains. Table 1 displays the complete set of models and datasets.
A.4. Installation
NotebookOS is deployed on AWS EC2 instances using provided Ansible playbooks (in setup/ansible [78]). After provisioning VMs with IAM roles granting S3 access (see the AWS documentation), the playbooks automate dependency installation, patching, and experiment orchestration across the cluster. The source code, as well as additional documentation, is available in the following GitHub repositories:
  1. (1)
  2. (2)
A.4.1. Multi-Node Deployments
The recommended way to deploy NotebookOS across multiple AWS EC2 virtual machines is using the provided Ansible playbooks (located in the setup/ansible directory). Before running any playbooks, there are a few configuration-related steps that must be performed. First, create a file called “all.yaml” in the setup/ansible/group_vars directory. The “all.template.yaml” file is provided as a starting point. There are five configuration parameters that must be specified explicitly:
  • ansible_ssh_private_key_file: Path to a private SSH key on the computer used to run the Ansible playbook(s). Used to enable access to the other VMs in the NotebookOS cluster.
  • private_key_to_upload: Path to a private SSH key on the computer used to run the Ansible playbook(s). This SSH key will be uploaded to the VMs in the NotebookOS cluster to enable SSH connectivity between them. This is useful because you may want to run some scripts or Ansible playbooks from one of the VMs once NotebookOS has been deployed.
  • public_key_to_upload: Path to a public SSH key on the computer used to run the Ansible playbook(s). This SSH key will be uploaded to the VMs in the NotebookOS cluster to enable SSH connectivity between them. This is useful because you may want to run some scripts or Ansible playbooks from one of the VMs once NotebookOS has been deployed.
  • gitbranch: the branch of the NotebookOS GitHub repository to use when deploying NotebookOS. You can default to using “main”.
  • git_personal_access_token: GitHub personal access token (PAT) with read access to the NotebookOS GitHub repository (or the fork of the NotebookOS source code that is being used).
To use Docker images built manually (rather than the provided images), there are several configuration parameters that must be changed. Please consult the documentation in the GitHub repository for additional details.
Once the “all.yaml” file has been created in the correct directory (i.e., “setup/ansible/group_vars”), you can begin deploying NotebookOS. First, run the playbook to create the Docker Swarm cluster:
\$ ansible-playbook -i inventory_file.ini create\_docker\_swarm\_cluster.yaml --tags ‘‘swarm’’
Next, deploy the Traefik [99] Docker Stack onto the Docker Swarm cluster. Traefik is an open source reverse proxy and ingress controller that NotebookOS uses to route external web traffic to the appropriate internal component. Traefik can be deployed onto the Docker Swarm cluster by executing the following command:
\$ ansible-playbook -i inventory_file.ini redeploy\_traefik\_docker\_stack.yaml
Finally, deploy the NotebookOS Docker Stack onto the Docker Swarm cluster:
\$ ansible-playbook -i inventory_file.ini deploy_distributed_notebook\_docker\_stack.yaml
If desired, verbose Ansible logging can be enabled by setting the ANSIBLE_STDOUT_CALLBACK environment variable to “debug”. For example:
\$ ANSIBLE_STDOUT_CALLBACK=debug ansible-playbook -i inventory_file.ini deploy\_distributed\_notebook\_docker\_stack.yaml
A.4.2. Single-Node and Development Deployments
For development, NotebookOS supports Docker Compose [22]. Run setup/install.sh to install host dependencies. A template for the Docker Compose yml file is provided in the deploy/docker-WSL2/ directory; generate it via the “generate-docker-compose-file.sh” script. Helper scripts and usage details are documented in the directory’s README. Once the docker-compose.yml file is generated, you can deploy NotebookOS via Docker Compose using the following command:
\$ docker compose up -d --build --scale daemon=4
In order for NotebookOS to operate correctly with 3 replicas per distributed kernel, the minimum value of the “--scale daemon=” argument is 3; however, the recommended minimum is 4 to enable kernel replica migrations to occur.
When deploying NotebookOS for development, we recommend specifying a few additional configuration parameters. First, the system expects that the core dumps will be written to a /cores directory. We recommend mounting a Docker volume so that data written to the /cores directory are persisted and can be accessed later. We also recommend deploying Dozzle [24] for easier log monitoring.
\$ docker run --name dozzle -d \
-v /var/run/docker.sock:/var/run/docker.sock:ro \
-p 7744:$80$ amir20/dozzle:latest
A.4.3. Running the Dashboard Locally or Independently
After cloning the dashboard repository, execute the following command from within the driver-frontend directory:
\$ npm install && npm run start:dev
Note that unless the above command is modified, the front-end server will become the active process for the terminal.
Next, to run the backend server, execute the following command from within the driver-backend directory:
\$ make run-server
Note that unless the avove command is modified, the back-end server will become the active process for the terminal.
A.4.4. AWS Configuration
To deploy NotebookOS on AWS, designate one internet-accessible EC2 VM with a public IPv4 address as the “primary VM”. This VM, which should be named ‘‘Jupyter NaaS Leader’’, hosts NotebookOS’s Traefik service, which routes external traffic (e.g., to the workload dashboard). Use the generate_inventory_file_aws.py file in the setup/ansible/ directory to generate an inventory.ini from inventory_file.template.ini. The script requires the primary VM to be named as above, and all other VMs in the NotebookOS cluster should be tagged with key ‘‘swarm’’ and value ‘‘follower’’. It assumes Ubuntu VMs with the ‘‘ubuntu’’ root username.
A.5. Experiment Workflow
The experiments for NotebookOS and the other baselines are primarily orchestrated using NotebookOS’s administrative dashboard [77]. The dashboard provides a browser-based interface to view and manage NotebookOS itself, as well as workloads running on NotebookOS.
A.6. Evaluation and Expected Results
The results presented in Section 5 correspond to the execution of a real-world workload trace on NotebookOS. We are working to open-source our workload trace. Executing the same workload on NotebookOS multiple times will generate approximately the same results, with small differences resulting from scheduling decisions and other random factors.
Appendix B Limited GPU Resource Availability
The demand for GPUs has surged in recent years, driven by the rapid growth of machine learning, deep learning, and high-performance computing applications [90, 60]. However, cloud providers such as Amazon Web Services (AWS), Google Cloud, and Microsoft Azure are struggling to meet this increasing demand due to limited GPU resources [108, 97, 21, 60]. This scarcity is further compounded by global supply chain challenges and the growing competition from sectors like gaming, cryptocurrency mining, and scientific research, all of which intensify the pressure on cloud infrastructure providers to allocate GPU resources efficiently [37, 98, 60]. The low availability of GPUs serves as further motivation of NotebookOS.
Appendix C Implementation
@@
Appendix D Simulation Study: Resource Usage
@@
Figure 14. Simulated GPU usage timeline over a 90-day snippet of the AdobeTrace spanning June through August.: Simulated GPU usage timeline over a 90-day snippet of the AdobeTrace spanning June through August.
@@
Appendix E Detailed Latency & Overhead Breakdown
Figure 15. Workflow steps of notebook cell execution requests in NotebookOS and other baselines.: Workflow steps of notebook cell execution requests in NotebookOS and other baselines.
Figure 16. Detailed end-to-end latency breakdown of execute requests messages processed by Reservation.: Detailed end-to-end latency breakdown of execute requests messages processed by Reservation.
Figure 17. Detailed end-to-end latency breakdown of execute requests messages processed by Batch.: Detailed end-to-end latency breakdown of execute requests messages processed by Batch.
Figure 18. Detailed end-to-end latency breakdown of execute requests messages processed by NotebookOS.: Detailed end-to-end latency breakdown of execute requests messages processed by NotebookOS.
Figure 19. Detailed end-to-end latency breakdown of execute requests messages processed by NotebookOS (LCP).: Detailed end-to-end latency breakdown of execute requests messages processed by NotebookOS (LCP).
@@ Appendix E Detailed Latency & Overhead Breakdown
E.1. Active Sessions & Training Events
Figure 20. The number of active user-submitted trainings and active user sessions during the full summer portion (i.e., June, July, and August) of the Adobe research cluster workload trace.: The number of active user-submitted trainings and active user sessions during the full summer portion (i.e., June, July, and August) of the Adobe research cluster workload trace.
Figure 20 presents a timeline plot showing the number of active sessions during the “summer portion” of the Adobe research cluster workload trace. Specifically, the number of active sessions is plotted on the secondary (i.e., right) y-axis. The summer portion includes data from the entirety of the months of June, July, and August. Initially, there are 0 active user sessions. By the end of June, July, and August, there are 206, 312, and 397 active user sessions, respectively. For June, the mean and median number of active user sessions are 115 and 130, respectively. For July, the mean and median number of active user sessions are 233 and 234, respectively. Finally, for August, the mean and median number of active user sessions are 379 and 385, respectively. The maximum number of active sessions at any given time throughout the summer Adobe research cluster workload trace is 433.
Figure 20 also presents a timeline plot showing the number of user-submitted training events being processed during the “summer portion” of the Adobe research cluster workload trace. Specifically, the number of active sessions is plotted on the primary (i.e., left) y-axis. By the end of June, July, and August, there are 58, 83, and 73 active user-submitted trainings, respectively. The mean and median number of active user-submitted trainings for the entire summer are 67.63 and 68, respectively. For June, the mean and median number of active user-submitted trainings are 31 and 30, respectively. For July, the mean and median number of active user-submitted trainings are 65 and 67, respectively. Finally, for August, the mean and median number of active user-submitted trainings are 105 and 108, respectively. The maximum number of active user-submitted trainings at any given time throughout the summer AdobeTrace is 141.